OpenStudy (anonymous):

use mathematical induction to prove that the statement is true for every positive integer n..... 2 is a factor of n^2-n+2

OpenStudy (amistre64):

does it work for the basis statement?

OpenStudy (amistre64):

since 1 is the first positive integer, it has to work out for n=1 1-1+2 = 2 and 2|2 is true. so now we have a basis to start from let n=k and rewrite as: k^2 - k + 2; now if its ture for some k, then is it true for some k+1? try it out (k+1)^2 - (k+1) + 2 ... does this simplify to the known format? (k^2+2k+1) - (k+1) + 2 k^2+2k+1 -k-1 + 2 k^2-k+2 + 2k but (k^2-k+2)|2 by hypothesis; and 2k|2 soo what can we conclude?

OpenStudy (anonymous):

so it is true

OpenStudy (amistre64):

its true without using induction as well

OpenStudy (amistre64):

\[\frac{n^2-n+2}{2}\] \[\frac{n(n-1)+2}{2}\] \[\frac{n(n-1)}{2}+\frac{2}{2}\] if n is odd, then n-1 is even and is divisible by 2 if n is even, then its divisible by 2 and that 2 is divisible by 2

OpenStudy (anonymous):

I really do appreciate your help!